OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the given point. y = x^3 − 2x + 1, (2, 5)

OpenStudy (el_tucan):

So i first solved for y' because that is you're slope or m in y - y1 = m(x - x1). y' = 3x^2 -2. plug in you're 2 to get 10. Plug in to the point slope formula.